# Understanding Multimodal LLMs Under Distribution Shifts: An Information-Theoretic Approach

### v1

@@ Abstract
Multimodal large language models (MLLMs) have shown promising capabilities but struggle under distribution shifts, where evaluation data differ from instruction tuning distributions. Although previous works have provided empirical evaluations, we argue that establishing a formal framework that can characterize and quantify the risk of MLLMs is necessary to ensure the safe and reliable application of MLLMs in the real world. By taking an information-theoretic perspective, we propose the first theoretical framework that enables the quantification of the maximum risk of MLLMs under distribution shifts. Central to our framework is the introduction of Effective Mutual Information (EMI), a principled metric that quantifies the relevance between input queries and model responses. We derive an upper bound for the EMI difference between in-distribution (ID) and out-of-distribution (OOD) data, connecting it to visual and textual distributional discrepancies. Extensive experiments on real benchmark datasets, spanning 61 shift scenarios empirically validate our theoretical insights.
@@ 1 Introduction
Multimodal large language models (MLLMs) have demonstrated remarkable capabilities in handling complex tasks that require reasoning over both visual and textual modalities. By leveraging visual instruction tuning (Liu et al., 2023; Dai et al., 2023; Zhu et al., 2024), MLLMs have shown promise in answering open-ended questions and generating contextually relevant captions. As a critical aspect for real-world deployment, MLLMs are expected to operate robustly in the wild, where distribution shifts occur—that is, when the evaluation data deviates from the instruction tuning data, whether due to changes in visual inputs (e.g, domain-specific images), textual inputs (e.g., linguistic variations), or the combination thereof. However, negative reports on MLLM failures under edge cases have steadily emerged, raising concerns about their reliability.
For example, MLLMs struggle with queries in specialized domains such as medical and chemistry (Zhang et al., 2024a; Han et al., 2024; Zhou et al., 2024), perform poorly on simple image classification tasks compared with open-ended question answering (Zhang et al., 2024b; Zhai et al., 2024), and frequently exhibit hallucination (Li et al., 2023b; Ye-Bin et al., 2025). Given the increasing impact of MLLMs, it is crucial to understand their failure modes under distribution shifts. Despite the significance of the problem, existing works often lack a fine-grained diagnosis for various factors of shifts. More importantly, the absence of a formal framework to explain the underlying principle further hinders the systematic understanding of MLLMs’ behavior. This motivates us to raise the research question:
@@
To address this, we propose an information-theoretic framework that characterizes MLLM performance under distribution shifts with theoretical rigor and practical interpretability. Our framework is well-suited for analyzing instruction-tuned models, which effectively maximizes the lower bound of mutual information between input query and response. Under this framework, we introduce effective mutual information (EMI) as a principled measurement for evaluating the relevance between an input query and model response. Intuitively, EMI is expected to be higher when a test input query originates from the in-distribution (ID) data similar to those used during instruction tuning, compared to when the input comes from out-of-distribution (OOD). To quantify this performance gap, we compute the difference in EMI between ID and OOD and derive an upper bound for this difference, expressed in terms of distributional discrepancies in input and output spaces (see Theorem 4.5 and 4.6). By grounding the measure in information theory, we provide the first theoretical framework to analyze and understand the impact of distribution shifts on MLLM performance.
Beyond the theoretical rigor, we further show that our framework holds practical value. In particular, we demonstrate that EMI is closely related to the widely used empirical evaluation metric for MLLMs, win rate (Ouyang et al., 2022) with an LLM judge (Zheng et al., 2023). The win rate commonly relies on external judge models such as GPT-4, thus it lacks mathematical guarantees due to the black-box nature of these evaluators. In contrast, EMI provides an theory-grounded measurement of the relevance between input queries and the output responses of the MLLM being evaluated, offering a principled metric for assessing performance under distribution shifts.
Finally, we conduct empirical validation for the theoretical framework and show that our theorems empirically hold in real-world benchmarks. Our experiments comprehensively examine 34 synthetic and 27 natural distribution shift scenarios, resulting in a total number of 61 ID-OOD evaluations for each MLLM. Results confirm strong correlations between EMI and win rate, as well as between EMI difference and its upper bounds, demonstrating the effectiveness of our framework in capturing performance gaps under diverse shifts. Our contributions can be summarized as follows:
  • We propose a new framework, effective mutual information (EMI), to analyze MLLM under distribution shift, and justify the use of EMI by showing the theoretical connection between EMI and win rate.
  • We derive a theoretical upper bound of MLLM’s performance gap which can be characterized by shifts over multimodal input queries and output discrepancies.
  • We empirically verify our theoretical statements on 61 real-world distribution shift scenarios of open-ended question-answering benchmarks with four MLLMs.
@@ 2 Preliminary
Let $\mathcal{X}=\mathcal{X}_{v}\times\mathcal{X}_{t}$ denote the input space, where $\mathcal{X}_{v}$ and $\mathcal{X}_{t}$ correspond to the visual and textual feature spaces, respectively. Similarly, let $\mathcal{Y}$ denote the response space. We define the random variables $\mathbf{X}=(X_{v},X_{t})\in\mathcal{X}$ and $Y\in\mathcal{Y}$, where $\mathbf{X}$ is the sequence of tokens that combine visual and text input queries, and $Y$ represents the associated response tokens. The joint population is denoted by $P_{\mathbf{X}Y}$, with marginals $P_{\mathbf{X}}$, $P_{Y}$, and the conditional distribution $P_{Y|\mathbf{X}}$. For subsequent sections, $P_{\mathbf{X}Y}$ refers to the instruction tuning data distribution which we consider as in-distribution (ID).
@@
MLLM usually consists of three components: (1) a visual encoder, (2) a vision-to-language projector, and (3) an LLM that processes a multimodal input sequence to generate a valid textual output $y$ in response to an input query $\mathbf{x}$. An MLLM can be regarded as modeling a conditional distribution $P_{\theta}(y|\mathbf{x})$, where $\theta$ is the model parameters. To attain the multimodal conversation capability, MLLMs commonly undergo a phase so-called visual instruction tuning (Liu et al., 2023; Dai et al., 2023) with an autoregressive objective as follows:
$\min_{\theta\in\Theta}\mathbb{E}_{\mathbf{x},y\sim P_{\mathbf{X}Y}}[\sum_{l=0}^{L}-\log P_{\theta}(y_{l}|\mathbf{x},y_{<l})],$ (1)
where $L$ is a sequence length and $y=(y_{0},...,y_{L})$. After being trained by Eq. (1), MLLM produces a response given a query of any possible tasks represented by text.
@@
(M)LLM-as-a-judge method (Zheng et al., 2023; Kim et al., 2024) is commonly adopted to evaluate open-ended generation. In this paradigm, a judge model produces preference scores or rankings for the responses given a query, model responses, and a scoring rubric. Among the evaluation metrics, the win rate (Eq. (2)) is one of the most widely used and representative.
Definition 2.1 (Win Rate).
Given a parametric reward function $r:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the win rate (WR) of model $P_{\theta}$ w.r.t. $P_{\mathbf{X}Y}$ are defined as follows:
$\begin{split}&\text{WR}(P_{\mathbf{X}Y};\theta):=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[\mathbb{I}(r(\mathbf{x},\hat{y})>r(\mathbf{x},y))],\end{split}$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function.
Here, the reward function $r(\cdot,\cdot)$, can be any possible (multimodal) LLMs such as GPT-4o (Hurst et al., 2024).
@@ 3 Motivation
While instruction-following MLLMs are designed to handle a diverse range of tasks, they often struggle with specialized domains (Zhang et al., 2024a; Zhou et al., 2024), perform poorly on simple image classification tasks (Zhai et al., 2024; Zhang et al., 2024b), and exhibit hallucinations (Li et al., 2023b; Ye-Bin et al., 2025). We argue that the fundamental cause of these failure modes in MLLMs can be traced back to distribution shifts. Specifically, the poor performance on classification tasks and specialized distributions can be attributed to shifts between instruction tuning distribution $P_{\mathbf{X}Y}$ and evaluation distribution $Q_{\mathbf{X}Y}$. Throughout this work, we comprehensively analyze three types of distributional shifts that can arise in MLLM:
@@
Although there have been similar observations of performance degradation of MLLM under distribution shifts (Achiam et al., 2023; Zhang et al., 2024a; Zhou et al., 2024; Zhang et al., 2024b), all of them present only the coarse empirical evaluation results without finer analysis on the underlying factor of those performance degradations. To the best of our knowledge, there is no formal framework to explain the performance variations of MLLMs in terms of distribution shifts— despite its crucial importance for ensuring reliable applications of MLLMs. To bridge this gap, we propose the first theoretical framework that characterizes MLLM performance variations under distribution shifts from an information-theoretic perspective.
@@ 4.1 Mutual Information for MLLM
A fundamental capability of MLLMs is their instruction-following property (Ouyang et al., 2022)—a direct outcome of instruction-tuning, where the model is trained to generate responses that are aligned with the intent of a given input query or instruction. To evaluate instruction-following, we first consider the mutual information (Shannon, 1948) to measure the shared information between the query and the corresponding model response.
@@
MI is deeply related to the entropy, which is defined as $H(P_{\mathbf{X}}):=-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[\log P_{\mathbf{X}}(\mathbf{x})]$. It is easy to check that $I(P_{\mathbf{X}Y})=H(P_{Y})-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{Y|\mathbf{X}=\mathbf{x}})]$. Intuitively, MI captures how much the response tells us about the query. One reason for considering MI is that the autoregressive objective in Eq. (1) effectively estimates MI (see Eq. (4.1)).
Visual instruction tuning effectively maximizes a lower bound of MI.
In Eq. (4.1), we show that the autoregressive objective for instruction tuning (Eq. (1)) effectively estimates the lower bound of MI when the model’s representation capacity is sufficiently high, i.e., when $\delta$ becomes small. Therefore, instruction tuning with autoregressive objective effectively maximize a lower bound of MI between the input query and desired response,
$I(P_{\mathbf{X}Y})=$ $\mathbb{E}_{\mathbf{x},y\sim P_{\mathbf{X}Y}}[\log P_{{\theta}}(y|\mathbf{x})]+H(P_{Y})+\delta$
@@
While $I(P_{\mathbf{X}Y})$ measures the MI between the input query and ground truth response from $P_{\mathbf{X}Y}$, one can measure MI between the input query and model response on the evaluation-time distribution. In particular, we use a tensor product $P_{\mathbf{X}}\otimes P_{\theta}$ to present this joint distribution between the input distribution $P_{\mathbf{X}}$ and generated output distribution $P_{\theta}(y|\mathbf{x})$:
@@
The challenge lies in the inherent variability of MI scales across data domains. Recall the formulation of $I(P_{\mathbf{X}}\otimes P_{\theta})$ in Eq. (6), the first term $H(\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[P_{\theta}(\cdot|\mathbf{x})])$ represents the upper bound of MI and varies with the data domain; and the second term $\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]$ reflects the input-output dependency and depends on the true data-generating process. For instance, given a fixed vocabulary, the responses from MLLM could contain more diverse words in a general domain whereas a narrower subset of words could be expected in the specialized domains such as medical. Then, $H(\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[P_{\theta}(\cdot|\mathbf{x})])$ and $\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]$ would be larger in a general domain. Therefore, we argue that a desired evaluation metric should disentangle the pure input-response relevance from the intrinsic characteristics of dataset.
@@ 4.2 Effective Mutual Information for Reliable MLLM Evaluation
Given the joint distribution $P_{\mathbf{X}Y}$ and MLLM $P_{\theta}$ parameterized with $\theta$, the effective mutual information between the input and model response is defined as below,
$\text{EMI}(P_{\mathbf{X}Y};\theta):=I(P_{\mathbf{X}}\otimes P_{\theta})-I(P_{\mathbf{X}Y}).$ (7)
Compared to the standard MI, $\text{EMI}(P_{\mathbf{X}Y};\theta)$ measures the “effective” relevance between the query $\mathbf{x}$ and the model response $\hat{y}$ by subtracting a ground truth MI $I(P_{\mathbf{X}Y})$ from $I(P_{\mathbf{X}}\otimes P_{\theta})$. Refer to Figure 5 in Appendix A, for an intuitive example: by accounting for a baseline level of MI, EMI quantifies the extent to which the model captures the effective relevance between the input and output. The use of EMI as an evaluation metric for MLLMs can be further supported by (1) its analogy to the excess risk and effective robustness; and (2) its connection to win rate.
@@
Connection to win rate.
We also show that EMI is closely related to win rate (i.e., Eq. (2)), a common metric used for evaluating MLLM generations. Conceptually, EMI quantifies the effective relevance between the input query and the model’s response by accounting for the baseline information, while win rate measures the preference of the model’s responses over a reference response. More formally, their connection can be mathematically established through the lens of a logit Bradley-Terry preference model (PM) formulation (Bradley & Terry, 1952; Hunter, 2004), $\text{logit}P(\hat{y}\succ y|\mathbf{x})$, a smooth and differentiable proxy for the discrete win rate function. To be specific, we commonly use $\log P(\hat{y}\succ y|\mathbf{x})$ to train a reward model (RM) which is adopted to compute the win rate (Eq. (2)). We compare both terms as below.
$\small\begin{split}{\text{PM}}(P_{\mathbf{X}Y};\theta):&=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[\text{logit}\;P(\hat{y}\succ y|\mathbf{x})]\\ &=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[r(\mathbf{x},\hat{y})-r(\mathbf{x},y)],\\ {\text{RM}}(P_{\mathbf{X}Y};\theta):&=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[\log\;P(\hat{y}\succ y|\mathbf{x})]\\ &=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[\log\sigma(r(\mathbf{x},\hat{y})-r(\mathbf{x},y))],\end{split}$
@@
${\text{PM}}(P_{\mathbf{X}Y};\theta)={\text{RM}}(P_{\mathbf{X}Y};\theta)-\log(1-e^{{\text{RM}}(P_{\mathbf{X}Y};\theta)}),$
${\text{RM}}(P_{\mathbf{X}Y};\theta)={\text{PM}}(P_{\mathbf{X}Y};\theta)-\log(1+e^{{\text{PM}}(P_{\mathbf{X}Y};\theta)}).$
Therefore, ${\text{PM}}(P_{\mathbf{X}Y};\theta)$ and ${\text{RM}}(P_{\mathbf{X}Y};\theta)$ exhibit a mutual equivalence, i.e., increase in ${\text{PM}}(P_{\mathbf{X}Y};\theta)$ corresponds to the increase in ${\text{RM}}(P_{\mathbf{X}Y};\theta)$, and vice versa. In Lemma 4.3, we establish an upper bound for the absolute difference between EMI and ${\text{PM}}(P_{\mathbf{X}Y};\theta)$, thereby demonstrating their closeness while ultimately highlighting the connection between EMI and win rate.
@@
$|\text{EMI}(P_{\mathbf{X}Y};\theta)-\text{PM}(P_{\mathbf{X}Y};\theta)|\leq\delta+4.4\delta^{\frac{1}{8}}.$
@@
Although we confine our analysis to $I(P_{\mathbf{X}Y})$ and $\text{EMI}(P_{\mathbf{X}Y};\theta)$, the chain rule, $I(P_{\mathbf{X}Y})=I(P_{X_{v}Y})+I(P_{X_{t}Y|X_{v}})$ allows EMI to further factorize the query-response relevance into two input modalities, which is suitable for multimodal LLM’s fine-grained evaluation.

### v2

@@ Abstract
[…]
Changdae Ohchangdae@cs.wisc.edu
@@ 1 Introduction
Multimodal large language models (MLLMs) have demonstrated remarkable capabilities in handling complex tasks that require reasoning over both visual and textual modalities. By leveraging visual instruction tuning (Liu et al., 2023; Dai et al., 2023; Zhu et al., 2024), MLLMs have shown promise in answering open-ended questions and generating contextually relevant captions. As a critical aspect for real-world deployment, MLLMs are expected to operate robustly in the wild, where distribution shifts occur—that is, when the evaluation data deviates from the instruction tuning data, whether due to changes in visual inputs (e.g, domain-specific images), textual inputs (e.g., linguistic variations), or the combination thereof. However, negative reports on MLLM failures in edge cases have steadily emerged, raising concerns about their reliability.
For example, MLLMs struggle with queries in specialized domains such as medicine and chemistry (Zhang et al., 2024a; Han et al., 2024; Zhou et al., 2024), perform poorly on simple image classification tasks compared to open-ended question answering (Zhang et al., 2024b; Zhai et al., 2024), and exhibit hallucinations to biased queries (Li et al., 2023b; Ye-Bin et al., 2025). Given the increasing impact of MLLMs, it is crucial to understand their failure modes under distribution shifts. Despite the significance of the problem, existing works often lack a fine-grained diagnosis for various factors of shifts. More importantly, the absence of a formal framework to explain the underlying principle further hinders the systematic understanding of an MLLM’s behavior. This motivates us to raise the research question:
@@
To address this, we propose an information-theoretic framework that characterizes MLLM performance under distribution shifts with theoretical rigor and practical interpretability. Our framework is well-suited for analyzing instruction-tuned models, wherein the purpose of learning is naturally connected to the mutual information between the input query and response. In this framework, we introduce effective mutual information (EMI) as a principled measure to access the relevance between an input query and model response. Intuitively, EMI is expected to be higher when a test input query originates from the in-distribution (ID) data, similar to those used during instruction tuning, compared to when the input comes from out-of-distribution (OOD). To quantify this performance gap, we compute the difference in EMI between ID and OOD and derive an upper bound for this difference, expressed in terms of distributional discrepancies in input and output spaces (see Theorem 4.5 and 4.6). By grounding the measure in information theory, we provide the first theoretical framework to analyze and understand the impact of distribution shifts on MLLM performance.
Beyond the theoretical rigor, we further show that our framework holds practical value. In particular, we demonstrate that EMI is closely related to a widely used empirical metric for MLLMs, relative preference score under the LLM-as-a-judge paradigm (Zheng et al., 2023). The relative preference score commonly relies on an external judge model, e.g., GPT-4o (Hurst et al., 2024), thus it lacks mathematical guarantees due to the black-box nature of these models. In contrast, EMI provides an theory-grounded measurement of the relevance between input queries and output responses of the MLLM being evaluated, offering a fundamental basis for assessing the MLLM’s performance gap under shifts.
Finally, we conduct comprehensive validation for the theoretical framework and show that our theorems empirically hold in real-world benchmarks. Our experiments comprehensively examine 34 synthetic and 27 natural distribution shift scenarios, resulting in a total number of 61 ID-OOD evaluations for each MLLM. Results confirm strong correlations between empirical estimates of EMI and relative preference, as well as correlations between the EMI difference and its upper bounds, demonstrating the effectiveness of our framework in capturing performance gaps under diverse shifts. Our contributions can be summarized as follows:
  • We propose a new framework, effective mutual information (EMI), to analyze MLLM under distribution shift, and justify the use of EMI by showing the theoretical connection between EMI and the LLM-judge-driven relative preference score.
  • We derive theoretical upper bounds of MLLM’s performance gap, which can be characterized by shifts over multimodal input queries and output discrepancies.
  • We empirically verify our theoretical statements on 61 real-world distribution shift scenarios of open-ended question-answering benchmarks with six MLLMs.
@@ 2 Preliminary
Let $\mathcal{X}=\mathcal{X}_{v}\times\mathcal{X}_{t}$ denote the input space, where $\mathcal{X}_{v}$ and $\mathcal{X}_{t}$ correspond to the visual and textual feature spaces, respectively. Similarly, let $\mathcal{Y}$ denote the response space. We define the random variables $\mathbf{X}=(X_{v},X_{t})\in\mathcal{X}$ and $Y\in\mathcal{Y}$, where $\mathbf{X}$ is the sequence of tokens that combine visual and text input queries, and $Y$ represents the associated response tokens. The joint distribution is denoted by $P_{\mathbf{X}Y}$, with marginals $P_{\mathbf{X}}$, $P_{Y}$, and the conditional distribution $P_{Y|\mathbf{X}}$. For subsequent sections, $P_{\mathbf{X}Y}$ refers to the instruction tuning data distribution, which we consider as in-distribution (ID).
@@
MLLM usually consists of three components: (1) a visual encoder, (2) a vision-to-language projector, and (3) an LLM backbone that processes a multimodal input sequence to generate a valid textual output $y$ in response to an input query $\mathbf{x}$. An MLLM can be regarded as modeling a conditional distribution $P_{\theta}(y|\mathbf{x})$, where $\theta$ is the model parameters. To attain the multimodal conversation capability, MLLMs commonly undergo a phase, so-called visual instruction tuning (Liu et al., 2023), with a conditional language modeling loss:
$\arg\min_{\theta\in\Theta}\mathbb{E}_{\mathbf{x},y\sim P_{\mathbf{X}Y}}[\sum_{l=1}^{L}-\log P_{\theta}(y_{l}|\mathbf{x},y_{<l})],$ (1)
where $L$ is a sequence length and $y=(y_{1},...,y_{L})$. After being trained by Eq. (1), MLLM produces a response given a query of any possible tasks represented by text.
@@
(M)LLM-as-a-judge method (Ouyang et al., 2022; Zheng et al., 2023; Kim et al., 2024) is commonly adopted to evaluate open-ended generation. In this paradigm, a judge model produces preference scores or rankings for the responses given a query, model responses, and a scoring rubric. Among the evaluation metrics, the relative preference score (RP score; Eq. (2)) is one of the most representative ones.
Definition 2.1 (Relative Preference Score).
Given a reward function $r:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the relative preference (RP) score of model $P_{\theta}$ w.r.t. $P_{\mathbf{X}Y}$ are defined as follows:
$\begin{split}&\text{RP}(P_{\mathbf{X}Y};P_{\theta}):=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[r(\mathbf{x},\hat{y})/r(\mathbf{x},y)].\end{split}$ (2)
Here, the reward function $r$ can be any possible MLLMs, such as GPT-4o (Hurst et al., 2024) or discriminative language models (Lambert et al., 2024), and we often take an output from another MLLM (usually more powerful than $P_{\theta}$) as a reference answer $y$.
@@ 3 Motivation
While instruction-following MLLMs are designed to handle a diverse range of tasks, they often struggle with specialized domains (Zhang et al., 2024a; Zhou et al., 2024), perform poorly on simple image classification tasks (Zhai et al., 2024; Zhang et al., 2024b), and hallucinate to biased queries (Li et al., 2023b; Ye-Bin et al., 2025). We argue that the fundamental cause of these failure modes in MLLMs can be traced back to distribution shifts. Specifically, poor performance on classification tasks and specialized distributions can be attributed to shifts between instruction tuning distribution $P_{\mathbf{X}Y}$ and evaluation distribution $Q_{\mathbf{X}Y}$. This work comprehensively analyzes three types of distributional shifts that can arise in MLLM as follow:
@@
Although there have been similar observations on the performance degradation of MLLM under distribution shifts (Achiam et al., 2023; Zhang et al., 2024a; Zhou et al., 2024; Zhang et al., 2024b), all of them present only the coarse empirical evaluation results without finer analysis on the underlying factor of those performance degradations. To the best of our knowledge, there is no formal framework to explain the performance variations of MLLMs in terms of distribution shifts— despite its crucial importance for ensuring reliable applications of MLLMs. To bridge this gap, we propose the first theoretical framework that characterizes MLLM performance variations under distribution shifts from an information-theoretic perspective.
@@ 4.1 Mutual Information for MLLM
A fundamental capability of MLLMs is their instruction-following property (Ouyang et al., 2022)—a direct outcome of instruction-tuning, where the model is trained to generate responses that are aligned with the intent of a given input query or instruction. To evaluate instruction-following capability, we first consider the mutual information (Shannon, 1948) to measure the shared information between the query and the corresponding model response.
@@
MI is deeply related to the entropy, which is defined as $H(P_{\mathbf{X}}):=-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[\log P_{\mathbf{X}}(\mathbf{x})]$. It is easy to check that $I(P_{\mathbf{X}Y})=H(P_{Y})-\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{Y|\mathbf{X}=\mathbf{x}})]$. Intuitively, MI captures how much the response tells us about the query and vice versa. One reason for considering MI for model evaluation is that the conditional language modeling objective in Eq. (1) is closely connected to MI as shown in Eq. (4.1).
Information-theoretic interpretation of instruction tuning.
In Eq. (4.1), we show that the objective for instruction tuning (Eq. 1) forms a lower bound of the MI between $\mathbf{X}$ and $Y$ subtracted by entropy over $Y$, given a sufficiently large representation capacity of the model, i.e., small $\delta$. As we are interested in measuring the input-output dependency, we focus on the $I(P_{\mathbf{X}Y})$ term as a metric to effectively gauge the upper bound of visual instruction tuning objective, Eq. (1).
$I(P_{\mathbf{X}Y})-H(P_{Y})=$ $\mathbb{E}_{\mathbf{x},y\sim P_{\mathbf{X}Y}}[\log P_{{\theta}}(y|\mathbf{x})]+\delta$
@@
While $I(P_{\mathbf{X}Y})$ measures the MI between the input query and ground truth response from $P_{\mathbf{X}Y}$, one may be interested in measuring MI between the query and model response on inference-time distributions. We use a tensor product $P_{\mathbf{X}}\otimes P_{\theta}$ to present a joint distribution between the input distribution $P_{\mathbf{X}}$ and model output distribution $P_{\theta}(y|\mathbf{x})$:
@@
The challenge lies in the inherent variability of MI scales across data domains. Recall the formulation of $I(P_{\mathbf{X}}\otimes P_{\theta})$ in Eq. (6), the first term $H(\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[P_{\theta}(\cdot|\mathbf{x})])$ represents the upper bound of MI and varies with the data domain; and the second term $\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]$ reflects the input-output dependency and depends on the true data-generating process. For instance, given a fixed vocabulary, the responses from MLLM could contain more diverse words in a general domain, whereas a narrower subset of words could be expected in the specialized domains, such as medical. Then, $H(\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[P_{\theta}(\cdot|\mathbf{x})])$ and $\mathbb{E}_{\mathbf{x}\sim P_{\mathbf{X}}}[H(P_{\theta}(\cdot|\mathbf{x}))]$ would be larger in a general domain. Therefore, we argue that a desired evaluation metric should disentangle the pure input-response relevance from the intrinsic characteristics of the dataset.
@@ 4.2 Effective Mutual Information for Reliable MLLM Evaluation
Given a joint distribution $P_{\mathbf{X}Y}$ and an MLLM $P_{\theta}$ parameterized with $\theta$, the effective mutual information between the input and model response is defined as below,
$\text{EMI}(P_{\mathbf{X}Y};P_{\theta}):=I(P_{\mathbf{X}}\otimes P_{\theta})-I(P_{\mathbf{X}Y}).$ (7)
Compared to the standard MI, $\text{EMI}(P_{\mathbf{X}Y};P_{\theta})$ measures the “effective” relevance between the query $\mathbf{x}$ and the model response $\hat{y}$ by subtracting a ground truth MI $I(P_{\mathbf{X}Y})$ from $I(P_{\mathbf{X}}\otimes P_{\theta})$. Refer to Figure 5 in Appendix A, for an intuitive example: by accounting for a baseline level of MI, EMI quantifies the extent to which the model captures the effective relevance between the input and output. The use of EMI as an evaluation metric for MLLMs can be further supported by (1) its analogy to the excess risk and effective robustness; (2) its connection to an LLM-judge score.
@@
Connection to LLM-judge score.
We also show that EMI is closely related to an LLM-judge-based metric, i.e., RP score in Eq. (2), a common metric used to assess MLLM outputs. Conceptually, EMI quantifies the effective relevance between the input query and the model’s response by accounting for the baseline mutual information, whereas the RP score measures the relative preference of the model’s responses over a reference response. More formally, their connection can be mathematically established through the lens of a logit Bradley-Terry preference model (PM) formulation (Bradley & Terry, 1952; Hunter, 2004), $\text{logit}P(\hat{y}\succ y|\mathbf{x})$, an alternative formulation of relative preference. To be specific, we commonly use $\log P(\hat{y}\succ y|\mathbf{x})$ to train a reward model (RM) which is adopted to compute LLM-judge score, such as Eq. (2). We compare both terms in the following.
$\small\begin{split}{\text{PM}}(P_{\mathbf{X}Y};P_{\theta}):&=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[\text{logit}\;P(\hat{y}\succ y|\mathbf{x})]\\ &=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[r(\mathbf{x},\hat{y})-r(\mathbf{x},y)],\\ {\text{RM}}(P_{\mathbf{X}Y};P_{\theta}):&=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[\log\;P(\hat{y}\succ y|\mathbf{x})]\\ &=\mathbb{E}_{\begin{subarray}{l}\mathbf{x},y\sim P_{\mathbf{X}Y}\\ \hat{y}\sim P_{\theta}(\cdot|\mathbf{x})\end{subarray}}[\log\sigma(r(\mathbf{x},\hat{y})-r(\mathbf{x},y))],\end{split}$
@@
${\text{PM}}(P_{\mathbf{X}Y};P_{\theta})={\text{RM}}(P_{\mathbf{X}Y};P_{\theta})-\log(1-e^{{\text{RM}}(P_{\mathbf{X}Y};P_{\theta})}),$
${\text{RM}}(P_{\mathbf{X}Y};P_{\theta})={\text{PM}}(P_{\mathbf{X}Y};P_{\theta})-\log(1+e^{{\text{PM}}(P_{\mathbf{X}Y};P_{\theta})}).$
Therefore, ${\text{PM}}(P_{\mathbf{X}Y};P_{\theta})$ and ${\text{RM}}(P_{\mathbf{X}Y};P_{\theta})$ exhibit a mutual equivalence, i.e., increase in ${\text{PM}}(P_{\mathbf{X}Y};P_{\theta})$ corresponds to the increase in ${\text{RM}}(P_{\mathbf{X}Y};P_{\theta})$, and vice versa. In Lemma 4.3, we establish an upper bound for the absolute difference between EMI and ${\text{PM}}(P_{\mathbf{X}Y};P_{\theta})$, thereby demonstrating their closeness and ultimately highlighting a connection between EMI and LLM-judge score, e.g., Eq (2).
@@
$|\text{EMI}(P_{\mathbf{X}Y};P_{\theta})-\text{PM}(P_{\mathbf{X}Y};P_{\theta})|\leq\delta+4.4\delta^{\frac{1}{8}}.$
@@
$\begin{split}&|\text{EMI}(P_{\mathbf{X}Y};P_{\theta}^{*})-\text{PM}(P_{\mathbf{X}Y};P_{\theta}^{*})|\leq\delta+4.4\delta^{\frac{1}{8}},\end{split}$
@@
Although we confine our analysis to $I(P_{\mathbf{X}Y})$ and $\text{EMI}(P_{\mathbf{X}Y};P_{\theta})$, the chain rule of MI, i.e., $I(P_{\mathbf{X}Y})=I(P_{X_{v}Y})+I(P_{X_{t}Y|X_{v}})$ allows us to further factorize the query-response relevance into two input modalities, which is suitable for multimodal LLM’s fine-grained evaluation.